# A Survey on Wi-Fi Sensing Generalizability: Taxonomy, Techniques, Datasets, and Future Research Prospects

### v2

@@ III Dataset
TABLE II: Open-sourced Wi-Fi Human Sensing Datasets
ID Dataset Main Domain Information Device Setting Highlight
Action Recognition
1 XRF55 [70] 4 environments, 39 participants, 55 actions 1Tx, 3Rx, 20MHz, 5GHz Multimodal, 55 actions
2 BullyDetect [211] 8 environments, 20 paired participants, 7 actions 1Tx, 1Rx, 20MHz, 5GHz Bullying actions
3 WiSDA [148] 1 environments, 3 participants, 6 actions 1Tx, 1Rx, 20MHz Widar3.0-like
4 WiNDR [63] 1 environments, 3 participants, 5 actions, 24 orientation 1Tx, 1Rx, 20MHz, 5GHz Full 360° coverage
5 WiMANS [283] 3 environments, 6 participants, 9 actions 1Tx, 1Rx, 20MHz, 2.4/5GHz Multi-person
6 WiGuesture [284] 1 environments, 8 participants, 6 actions 1Tx, 1Rx, 100Hz, 2.4GHz ESP32S3
7 RoGER [225] 2 environments, 4 participants, 6 actions, 4 orientation 1Tx, 2Rx, 20MHz, 2.437GHz AR9580
8 ImgFi [285] 1 environments, 5 participants, 6 actions 1Tx, 1Rx, 20MHz Converted to images
9 ResMon [224] 2 environments, 3 areas, 6 participants, 4 actions 1Tx, 1Rx, 20/40MHz, 2.4/5GHz Cross-band test
10 Meneghello et al. [286] 6 environments, 4 participants, 7 actions 1Tx, 1Rx, 80MHz, 5.21GHz 80MHz
11 Demrozi et al. [287] 2 environments, 6 participants, 6 actions 2Tx, 2Rx, 2.4GHz Nexmon firmware
12 MM-Fi [72] 4 environments, 40 participants, 27actions 1TX, 1RX, 20MHz, 5GHz Multimodal
13 NTU-Fi HAR [288] 1 environments, 20 participants, 6 actions 1Tx, 1Rx, 40MHz, 5GHz TP-Link N750
14 FallDar [289] 4 environments, 6 locations, 6 participants, fall & normal 1Tx, 1Rx, 1000Hz, 5GHz 6 months
15 ReWis [66] 3 environments, 2 participants, 4 actions 1Tx, 3Rx, 20/80MHz, 5GHz 80MHz
16 OPERAnet [290] 2 environments, 6 participants, 6 actions 1Tx, 2Rx, 5GHz Multimodal
17 SHARP [91] 3 environments, 4 locations, 3 participants, 7 actions 1Tx, 1Rx, 80MB, 5GHz 80MHz
18 CSI-HAR-Dataset [291] 1 environments, 3 participants, 7 actions 1Tx, 1Rx, 20MHz, 5GHz 802.11ac
19 CSIDA [193] 2 environments, 5 locations, 5 participants, 6 actions 1Tx, 1Rx, 40MHz, 5GHz 114 subcarriers
20 RISE [237] 2 environments, 6 participants, 6 actions 1Tx, 1Rx multi-rf devices
21 HTHI [292] 1 environment, 40 paired, 12 actions 2Tx, 3Rx, 20MHz, 2.4GHz Human-human actions
22 DeepSeg [261] 1 environment, 5 participants, 10 actions 1Tx, 3Rx fine-/coarse-grained actions
23 RF-Net [199] 6 environments, 11 participants, 6 actions 2 Tx-Rx pairs, 20MHz 120 environment
24 Baha’A et al. [293] 3 environments, 30 participants, 12 actions 1Tx, 3Rx, 20MHz, 2.4GHz LOS/NLOS
25 Widar 3.0 [67] 3 environments, 5 locations, 5 orientations, 16 participants, 16 actions 1Tx, 6Rx, 20MHz, 5.825GHz Rich domains
26 ARIL [15] 16 locations, 1 participants, 6 actions 1Tx, 1Rx, 20MHz, 2.4GHz USRP data, clean phase
27 WiAR [294] 3 environments, 10 participants, 16 actions 1Tx, 1Rx, 20MHz, 5GHz RSSI and CSI
28 SAR [295, 221] 1 environment, 9 participants, 6actions 2-3Tx, 3Rx, 20MHz, 2.4GHz 6 days
29 FallDeFi [39] 5 environments, 3 participants, 9 actions 2Tx, 2Rx, 20MHz, 5.2GHz Fall detection
30 SignFi [14] 2 environments, 5 participants, 276 gestures 3Tx, 1Rx, 20MHz, 5GHz 276 ASL Gestures
31 CrossSense [158] 3 environments, 15 locations, 100 participants, 40 actions 1Tx, 1Rx, 20MHz, 5GHz 100 participants
32 Yousef et al. [62] 1 environments, 6 participants, 6 actions 1Tx, 1Rx, 20MHz, 5GHz Fall
@@ 30 SignFi [14] 2 environments, 5 participants, 276 gestures 3Tx, 1Rx, 20MHz, 5GHz 276 ASL Gestures
33 NTU-Fi HumanID [191] 2 environments, 20 participants, walking 1Tx, 1Rx, 40MHz, 5GHz TP-Link N750
20 RISE [237] 3 environments, 15 participants 1Tx, 1Rx multi-rf devices
34 GaitID [30] 2 environments, 11 participants, 8 directions 1Tx, 6Rx 8 directions
31 CrossSense [158] 3 environments, 100 participants 1Tx, 1Rx, 20MHz, 5GHz 100 participants
@@
35 XRFv2 [296] 3 environments, 16 participants, 30 actions 1Tx, 3Rx, 20MHz, 5GHz Multimodal, continuous actions
36 Person-in-WiFi 3D [50] 3 environments, 7 participants, 8actions 1Tx, 3Rx, 20MHz, 5GHz Multi-person
1 XRF55 [70] 4 environments, 39 participants, 55 actions 1Tx, 3Rx, 20MHz, 5GHz Multimodal, 55 actions
12 MM-Fi [72] 4 environments, 40 participants, 27 actions 20MHz, 5GHz Multimodal
@@
37 MetaLoc [197] 2 environments, 180 locations 3Tx, 1Rx, 20MHz, 5GHz RSSI and CSI
38 Chen and Chang [297] 2 environments, 16 locations 2Tx, 2Rx CSI fingerprints
26 ARIL [15] 1 environments, 16 locations, 1 participants, 6 actions 1Tx, 1Rx, 20MHz, 2.4GHz USRP data, clean phase
@@ 38 Chen and Chang [297] 2 environments, 16 locations 2Tx, 2Rx CSI fingerprints
39 SDP [298] 4 scenarios multi-APs, 80/160MHz, 5GHz 0-2 people
40 DASECount [227] 2 environments, 2 scenarios (NLOS/LOS), 3 motions 1Tx, 1Rx, 40MHz, 2.4GHz 0-8 people
10 Meneghello et al. [286] 1 environments, 10 participants 1Tx, 1Rx, 80MHz, 5.21GHz 1-10 people
@@ 39 SDP [298] 4 scenarios multi-APs, 80/160MHz, 5GHz 0-2 people
35 XRFv2 [296] 3 environments, 16 participants, 30 actions 1Tx, 3Rx, 20MHz, 5GHz Multimodal, continuous actions
41 WiFiTAD [299] 1 environment, 3 participants, 7 actions 1Tx, 1Rx, 5GHz Continuous actions
@@
42 CSI-Bench [73] 26 environments, 35 participants 16 device configuration In-the-wild data, 461 hours
10 Meneghello et al. [286] 7 environments, 13 participants 1Tx, 1Rx, 80MHz, 5.21GHz Three tasks, 80MHz
43 CSI-Net [210] 1 environments, 5 positions, 30 participants 1Tx, 1Rx, 20MHz, 5GHz Four tasks, Biometrics
Figure 17: Bubble chart of public Wi-Fi sensing datasets, where the bubble size indicates the number of samples and the vertical axis represents the number of classes. For visual clarity, datasets with extreme values, such as SignFi [14] (296 classes) and CSI-Bench [73] (over 1,000,000 samples), are excluded from the plot. Similarly, datasets like WiFiTAD [299], MM-Fi [72], and XRFv2 [296], which collect continuous sequences over time without clear per-sample annotations, are also not included due to the ambiguity in estimating their sample counts.
@@ 42 CSI-Bench [73] 26 environments, 35 participants 16 device configuration In-the-wild data, 461 hours
(3) WiSDA [148] dataset was collected in an office setting, equipped with a desk and chair, within a 2m×$\times$2m sensing area. Three participants performed six activity categories: Push, Sweep, Clap, Slide, Draw-Z, and Draw-N. Each person collected 100 samples on 30 subcarriers and converted them to images, resulting in 3000 samples per action. The data was captured using a single pair of antennas, with one acting as the transmitter and the other as the receiver within the sensing area, similar to the setup used in the Widar 3.0 dataset. The bandwidth is 20 MHz.
(4) WiNDR [63] dataset was collected in an indoor office room measuring 4.8m×$\times$3.1m. Three volunteers incluing two males and one female, aged between 25 and 30 years old and weighing between 50 to 65 kg, were asked to perform five distinct hand gestures: drawing a circle (start/accept), crossing hands (stop), clapping (switch context/menu), raising hands (increase/turn up), and lowering hands (decrease/turn down). The volunteers were instructed to perform these gestures in 24 different directions, marked by gray lines on the floor, with each direction spaced 15 degrees apart. These gestures were performed 40 times in each of the 24 directions, sampled every 15 degrees around the subjects. The collection of gesture data was facilitated by two computers, the antennas were extended with 3.5m cables and positioned 1.2m above the ground using tripods.The sampling rate was 200Hz.
@@
(9) ResMon [224] dataset was collected in two environments: a laboratory and a meeting room, with the laboratory further collected at two areas. Six volunteers performed 4 actions: stable breath, cough, sneeze, and yawn. The experimental setup included a two-antenna transmitter and a three-antenna receiver, operating at 2.437 GHz with a 20MHz bandwidth and sampling rate of 100 Hz. ResMon also includes samples collected at 5GHz with 40MHz bandwidth for cross-band evaluation.
(10) Meneghello et al. [286] present a Wi-Fi sensing dataset collected using the 802.11ac protocol with an 80 MHz bandwidth centered at 5,210 MHz. It supports action recognition (AR), person identification (PI), and people counting (PC) tasks. For AR, four participants performed up to seven activities (e.g., walking, running, jumping, sitting still, standing still, sitting down/standing up, arm exercises) in six environments: including a bedroom, living room, kitchen, university laboratory, university office, and semi-anechoic chamber. For PI, 10 individuals were recorded moving freely in a meeting room (7 × 7.5 × 3.5 m). For PC, 1 to 10 people moved simultaneously in the meeting room. The dataset spans over 13 hours of channel state information (CSI) recordings, collected using the Nexmon tool with Netgear X4S AC2600, Asus RT-AC86U, and TP-Link AD7200 routers.
(11) Demrozi et al. [287] dataset was collected in two office environments. The first office, measuring 12m×$\times$6m×$\times$3m, is divided into two blocks of workspaces, each equipped with a Raspberry Pi device and an access point (AP). The second office is smaller, 6m×$\times$4m×$\times$2.75m, and also contains two Raspberry Pi devices and two APs. The Raspberry Pis and APs were paired to collect Channel State Information (CSI). Six participants, including two females and four males, performed five types of activities: entering the office, walking, standing, sitting, and leaving the office. The devices were arranged such that two wireless routers were placed 5 meters apart at a height of 140 cm, creating two separate 2.4 GHz WiFi networks. The total size of the dataset is approximately 70 GB.
@@
(14) FallDar [289] dataset, developed for fall detection, was collected in two environments: a home environment and an office environment. In the home environment, five participants performed fall actions at three locations—dining room, living room, and balcony—resulting in 115 fall samples. The transmitter and receiver were placed in two settings: line-of-sight (LoS) and non-line-of-sight (NLoS). In the office environment, the transmitter was fixed in position, while the receiver was placed at 16 locations sequentially. Two participants contributed to the dataset. One performed fall actions at locations 1 and 3, while the other performed falls at locations 1 and 4. Additionally, one participant walked freely in the office to collect normal activity data. This resulted in 228 fall samples and 504 normal activity samples. In both environments, the transmitter had one antenna, and the receiver had three antennas. Data was collected at a sampling rate of 1,000 Hz with a center frequency of 5.825 GHz.
@@
(24) Baha’A et al. [293] data collected in three different environments: two with line-of-sight (LOS) settings and one with a non-line-of-sight (NLOS) setting. Thirty volunteers performed five different actions: falling from a sitting position, falling from a standing position, walking, sitting down and standing up, and picking up a pen from the ground. Each action was performed 20 times. The experimental setup included a single-antenna transmitter and a three-antenna receiver, operating at a frequency of 2.4 GHz with a sampling rate of 320 Hz.
(25) Widar3.0 [67] dataset was collected in three environments: a classroom, a spacious hall, and an office room. In each environment, volunteers performed actions within a 2m $\times$ 2m area, with five designated locations and five orientations at each location. The dataset includes two types of hand gestures. The first type consists of general hand movements, including pushing and pulling, sweeping, clapping, sliding, drawing circles, and drawing zigzags. The second type involves drawing the numbers 0$\sim$9. The first type of action was performed by 16 participants, resulting in 12,000 samples (16$\times$5$\times$5$\times$6$\times$5 instances), while the second type was demonstrated by 2 participants, with 5,000 samples (2$\times$5$\times$5$\times$10$\times$10 instances). Widar3.0 was collected using a transmitter with one antenna and at least three receivers, each with three antennas. The dataset was captured on channel 165 at 5.825 GHz with a sampling rate of 1,000 Hz.
@@
(32) Yousef et.al. 2017 [62]dataset was collected in an indoor office area where the WiFi transmitter and receiver were positioned 3m apart in a line-of-sight (LOS) condition. Six volunteers performed six different activities: lying down, falling down, walking, running, sitting down, and standing up. Each activity was performed 20 times, yielding 720 samples, and each activity was recorded for 20 seconds. Additionally, video recordings were made during the experiments to help label and annotate the activities in the dataset. The sampling rate was 1000 Hz.
@@
(40) DASECount  [227] dataset, designed for crowd counting, was collected in two environments: an office and a lecture hall, each with both line-of-sight (LOS) and non-line-of-sight (NLOS) configurations, resulting in four different environmental settings. Volunteers, ranging from 0 to 8 people, performed three types of motion: static (seated with free actions like eating or typing), dynamic (random walking), and mixed (unrestricted activities including walking and sitting). Each sampling sequence was recorded for 5 minutes. The experimental setup included a two-antenna transmitter and a three-antenna receiver, operating at 2.4 GHz with a bandwidth of 40 MHz and a sampling rate of 100 Hz.

### v4

@@ 16 OPERAnet [15] 2 environments, 6 participants, 6 actions 1Tx, 2Rx, 5GHz Multimodal 71%-100%
20 RISE [270] 2 environments, 6 participants, 6 actions 1Tx, 1Rx multi-rf devices 93.9%-100%
21 HTHI [4] 1 environment, 40 paired, 12 actions 2Tx, 3Rx, 20MHz, 2.4GHz Human-human actions note reported
22 DeepSeg [243] 1 environment, 5 participants, 10 actions 1Tx, 3Rx fine-/coarse-grained actions 94%
23 RF-Net [42] 6 environments, 11 participants, 6 actions 2 Tx-Rx pairs, 20MHz 120 environment $\approx$ 80%
24 Baha’A et al. [8] 3 environments, 30 participants, 12 actions 1Tx, 3Rx, 20MHz, 2.4GHz LOS/NLOS not reported
25 Widar 3.0 [313] 3 environments, 5 locations, 5 orientations, 16 participants, 16 actions 1Tx, 6Rx, 20MHz, 5.825GHz Rich domains 92.7%-92.9%
26 ARIL [204] 16 locations, 1 participants, 6 actions 1Tx, 1Rx, 20MHz, 2.4GHz USRP data, clean phase 89.57%
27 WiAR [69] 3 environments, 10 participants, 16 actions 1Tx, 1Rx, 20MHz, 5GHz RSSI and CSI 90.62%-96.25%
28 SAR [16, 17] 1 environment, 9 participants, 6actions 2-3Tx, 3Rx, 20MHz, 2.4GHz 6 days 60%-100%
29 FallDeFi [156] 5 environments, 3 participants, 9 actions 2Tx, 2Rx, 20MHz, 5.2GHz Fall detection 80.10%-88.90%
30 SignFi [142] 2 environments, 5 participants, 276 gestures 3Tx, 1Rx, 20MHz, 5GHz 276 ASL Gestures 94.81%-98.91%
31 CrossSense [283] 3 environments, 15 locations, 100 participants, 40 actions 1Tx, 1Rx, 20MHz, 5GHz 100 participants 90%
32 Yousef et al. [262] 1 environments, 6 participants, 6 actions 1Tx, 1Rx, 20MHz, 5GHz Fall 81%-97%
[…]
@@ 22 DeepSeg [243] 1 environment, 5 participants, 10 actions 1Tx, 3Rx fine-/coarse-grained actions 94%
34 NTU-Fi HumanID [201] 2 environments, 20 participants, walking 1Tx, 1Rx, 40MHz, 5GHz TP-Link N750 86.29&-98.34%
20 RISE [270] 3 environments, 15 participants 1Tx, 1Rx multi-rf devices 97.6%
35 GaitID [299] 2 environments, 11 participants, 8 directions 1Tx, 6Rx 8 directions 61%-100%
31 CrossSense [283] 3 environments, 100 participants 1Tx, 1Rx, 20MHz, 5GHz 100 participants 80%
@@
[…]
36 XRFv2 [112] 3 environments, 16 participants, 30 actions 1Tx, 3Rx, 20MHz, 5GHz Multimodal, continuous actions not reported
37 Person-in-WiFi 3D [251] 3 environments, 7 participants, 8actions 1Tx, 3Rx, 20MHz, 5GHz Multi-person 107.2mm MPJPE
1 XRF55 [208] 4 environments, 39 participants, 55 actions 1Tx, 3Rx, 20MHz, 5GHz Multimodal, 55 actions not reported
12 MM-Fi [254] 4 environments, 40 participants, 27 actions 20MHz, 5GHz Multimodal 197.1mm MPJPE
@@
[…]
39 Chen and Chang [23] 2 environments, 16 locations 2Tx, 2Rx CSI fingerprints 49.47%-90.44%
26 ARIL [204] 1 environments, 16 locations, 1 participants, 6 actions 1Tx, 1Rx, 20MHz, 2.4GHz USRP data, clean phase 95.68%
@@
40 SDP [72] 4 scenarios multi-APs, 80/160MHz, 5GHz 0-2 people not reported
41 DASECount [83] 2 environments, 2 scenarios (NLOS/LOS), 3 motions 1Tx, 1Rx, 40MHz, 2.4GHz 0-8 people 93.29%-99.17%
10 Meneghello et al. [147] 1 environments, 10 participants 1Tx, 1Rx, 80MHz, 5.21GHz 1-10 people not reported
@@ 40 SDP [72] 4 scenarios multi-APs, 80/160MHz, 5GHz 0-2 people not reported
36 XRFv2 [112] 3 environments, 16 participants, 30 actions 1Tx, 3Rx, 20MHz, 5GHz Multimodal, continuous actions 78.74 mAP
41 WiFiTAD [139] 1 environment, 3 participants, 7 actions 1Tx, 1Rx, 5GHz Continuous actions 74.5 mAP
@@
43 CSI-Bench [323] 26 environments, 35 participants 16 device configuration In-the-wild data, 461 hours 94.88% fall detection, etc.
10 Meneghello et al. [147] 7 environments, 13 participants 1Tx, 1Rx, 80MHz, 5.21GHz Three tasks, 80MHz not reported
44 CSI-Net [206] 1 environments, 5 positions, 30 participants 1Tx, 1Rx, 20MHz, 5GHz Four tasks, Biometrics 96.67% fall detection, etc.
Figure 17: Bubble chart of public Wi-Fi sensing datasets, where the bubble size indicates the number of samples and the vertical axis represents the number of classes. For visual clarity, datasets with extreme values, such as SignFi [142] (296 classes) and CSI-Bench [323] (over 1,000,000 samples), are excluded from the plot. Besides, datasets like WiFiTAD [139], MM-Fi [254], and XRFv2 [112], which collect continuous sequences over time without clear per-sample annotations, are also not included due to the ambiguity in estimating their sample counts.
@@
(3) WiSDA [99] dataset was collected in an office setting, equipped with a desk and chair, within a $\times2m$ sensing area. Three participants performed six activity categories: Push, Sweep, Clap, Slide, Draw-Z, and Draw-N. Each person collected 100 samples on 30 subcarriers and converted them to images, resulting in 3000 samples per action. The data was captured using a single pair of antennas, with one acting as the transmitter and the other as the receiver within the sensing area, similar to the setup used in the Widar 3.0 dataset. The bandwidth is 20 MHz.
[…]
@@
(9) ResMon [310] dataset was collected in two environments: a laboratory and a meeting room, with the laboratory further collected at two areas. Six volunteers performed 4 actions: stable breath, cough, sneeze, and yawn. The experimental setup included a two-antenna transmitter and a three-antenna receiver, operating at 2.437 GHz with a 20MHz bandwidth and sampling rate of 100 Hz. ResMon also includes samples collected at 5 GHz with a 40 MHz bandwidth for cross-band evaluation.
(10) Meneghello et al. [147] present a Wi-Fi sensing dataset collected using the 802.11ac protocol with an 80 MHz bandwidth centered at 5,210 MHz. It supports action recognition (AR), person identification (PI), and people counting (PC) tasks. For AR, four participants performed up to seven activities (e.g., walking, running, jumping, sitting still, standing still, sitting down/standing up, arm exercises) in six environments: including a bedroom, living room, kitchen, university laboratory, university office, and semi-anechoic chamber. For PI, 10 individuals were recorded moving freely in a meeting room (7 × 7.5 × 3.5 m). For PC, 1 to 10 people moved simultaneously in the meeting room. The dataset spans over 13 hours of Channel State Information recordings, collected using the Nexmon tool with Netgear X4S AC2600, Asus RT-AC86U, and TP-Link AD7200 routers.
[…]
@@
(14) FallDar [258] dataset, developed for fall detection, was collected in two environments: a home environment and an office environment. In the home environment, five participants performed fall actions at three locations—dining room, living room, and balcony—resulting in 115 fall samples. The transmitter and receiver were placed in two settings: line-of-sight and non-line-of-sight. In the office environment, the transmitter was fixed in position, while the receiver was placed at 16 locations sequentially. Two participants contributed to the dataset. One performed fall actions at locations 1 and 3, while the other performed falls at locations 1 and 4. Additionally, one participant walked freely in the office to collect normal activity data. This resulted in 228 fall samples and 504 normal activity samples. In both environments, the transmitter had one antenna, and the receiver had three antennas. Data was collected at a sampling rate of 1,000 Hz with a center frequency of 5.825 GHz.
@@
(24) Baha’A et al. [8] data collected in three different environments: two with line-of-sight settings and one with a non-line-of-sight setting. Thirty volunteers performed five different actions: falling from a sitting position, falling from a standing position, walking, sitting down and standing up, and picking up a pen from the ground. Each action was performed 20 times. The experimental setup included a single-antenna transmitter and a three-antenna receiver, operating at a frequency of 2.4 GHz with a sampling rate of 320 Hz.
[…]
@@
(32) Yousef et.al. 2017 [262]dataset was collected in an indoor office area where the WiFi transmitter and receiver were positioned 3m apart in a line-of-sight condition. Six volunteers performed six different activities: lying down, falling down, walking, running, sitting down, and standing up. Each activity was performed 20 times, yielding 720 samples, and each activity was recorded for 20 seconds. Additionally, video recordings were made during the experiments to help label and annotate the activities in the dataset. The sampling rate was 1000 Hz.
(33) OctoNet [264] is collected in three indoor environments (office, laboratory, and living room) within a 4.1m $\times$ 4.1m area. It includes 41 participants performing 62 daily activities. Participants repeatedly performed each assigned activity following standardized instructions, including both structured routines and freestyle movements. For Wi-Fi sensing, one transmitter with a single antenna operates at 5.18 GHz with 40 MHz bandwidth. Four dual-antenna receivers are arranged in a rectangular layout, forming eight Wi-Fi links (1 Tx $\times$ 4 Rx). The CSI sampling rate is 75.62 Hz, yielding approximately 27.35 million synchronized frames. OctoNet further integrates 12 synchronized modalities and provides high-fidelity 3D skeletal keypoints captured by a motion-capture system, enabling research on multimodal learning, cross-modal tasks, and 3D human pose estimation.
@@
(41) DASECount  [83] dataset, designed for crowd counting, was collected in two environments: an office and a lecture hall, each with both line-of-sight and non-line-of-sight configurations, resulting in four different environmental settings. Volunteers, ranging from 0 to 8 people, performed three types of motion: static (seated with free actions like eating or typing), dynamic (random walking), and mixed (unrestricted activities including walking and sitting). Each sampling sequence was recorded for 5 minutes. The experimental setup included a two-antenna transmitter and a three-antenna receiver, operating at 2.4 GHz with a bandwidth of 40 MHz and a sampling rate of 100 Hz.